Mixture problem: how many liters of 10% solution and how many liters of 20% solution must be mixed to obtain 20 liters of a 17.5% solution? show equations and solve
Answer by nerdybill(7384)

Mixture problem: how many liters of 10% solution and how many liters of 20% solution must be mixed to obtain 20 liters of a 17.5% solution? show equations and solve
.
Let x = liters of 10% solution
then
20-x = liters of 20% solution
.
.10x + .20(20-x) = .175(20)
.10x + 4 - .20x = 3.5
4 - .10x = 3.5
-.10x = -.5
x = -.5/-.10
x = 5 liters (10% solution)
.
20% solution:
20-x = 20-5 = 15 liters